Determine the mass (in grams) of each of the following:

0.0872 mol acetylene, C2H2

Do not report in scientific notation.
Answer #1

Molar mass of C2H2,

MM = 2*MM(C) + 2*MM(H)

= 2*12.01 + 2*1.008

= 26.036 g/mol

mass of C2H2,

m = number of mol * molar mass

= 8.72*10^-2 mol * 26.036 g/mol

= 2.27 g

Answer: 2.27 g